# Automated Creation of Source Code Variants of a Cryptographic Hash Function Implementation Using Generative Pre-Trained Transformer Models

### v1

@@ Abstract
Generative pre-trained transformers (GPT’s) are a type of large language machine learning model that are unusually adept at producing novel, and coherent, natural language. Notably, these technologies have also been extended to computer programming languages with great success. However, GPT model outputs in general are stochastic and not always correct. For programming languages, the exact specification of the computer code, syntactically and algorithmically, is strictly required in order to ensure the security of computing systems and applications. Therefore, using GPT models to generate computer code poses an important security risk – while at the same time allowing for potential innovation in how computer code is generated. In this study the ability of GPT models to generate novel and correct versions, and notably very insecure versions, of implementations of the cryptographic hash function SHA-1 is examined. The GPT models Llama-2-70b-chat-hf, Mistral-7B-Instruct-v0.1, and zephyr-7b-alpha are used. The GPT models are prompted to re-write each function using a modified version of the localGPT framework and langchain to provide word embedding context of the full source code and header files to the model, resulting in over 130,000130000130,000130 , 000 function re-write GPT output text blocks (that are potentially correct source code), approximately 40,0004000040,00040 , 000 of which were able to be parsed as C code and subsequently compiled. The generated code is analyzed for being compilable, correctness of the algorithm, memory leaks, compiler optimization stability, and character distance to the reference implementation. Remarkably, several generated function variants have a high implementation security risk of being correct for some test vectors, but incorrect for other test vectors. Additionally, many function implementations were not correct to the reference algorithm of SHA-1, but produced hashes that have some of the basic characteristics of hash functions. Many of the function re-writes contained serious flaws such as memory leaks, integer overflows, out of bounds accesses, use of uninitialised values, and compiler optimization instability. Compiler optimization settings and SHA-256 hash checksums of the compiled binaries are used to cluster implementations that are equivalent but may not have identical syntax - using this clustering over 100,000100000100,000100 , 000 distinct, novel, and correct versions of the SHA-1 codebase were generated where each component C function of the reference implementation is different from the original code.
@@ 1 Introduction
In this study we explore whether current GPT models can be used to generate correct algorithmic invariant implementations of a cryptographic hash function in C code. Specifically, we examine the task of re-writing an implementation of the cryptographic hash function known as SHA-1 [SHA1_NIST_specifications]. An interesting byproduct of code-rewriting with the GPT models is that a large number of the implementations are wrong in a variety of surprising ways, including containing software risks. These include compiler optimization instability (meaning that the output changes based on what compiler optimization settings were used), memory leaks, integer overflows, out of bounds writes, and implementations that are correct for some test vectors but not correct for other test vectors. Implementation risks in cryptographic algorithms is a critically important type of bug that exists in cryptography library implementations [cryptoeprint:2023/331, 9426079, 6494288, 8405614]. This study serves to caution that arbitrarily using GPT models for creating, or re-writing, source code can introduce serious flaws. GPT models, with their current capabilities, work well as research tools to study interesting source code variants, but using them for practical code generation poses a software security risk.
Importantly, SHA-1 [SHA1_NIST_specifications] is considered broken since at least one method of generating a SHA-1 hash collision has been demonstrated [stevens2017first], and is generally considered not secure for validating the integrity of data based on a variety of attacks [wang2005finding, biham2005collisions, karpman2015practical, stevens2013new, 255248]. However, SHA-1 is still widely deployed in information technology systems, and moreover its implementation is relatively short compared to other cryptographic algorithms, making it a good candidate for a proof of concept study on the capabilities of using GPT models for re-writing a cryptographic algorithm implementation.
@@
Automated code re-writing, in particular of cryptographic algorithm implementations, is an excellent test case for evaluating GPT models capabilities involving computer code because the implementation must be exactly correct or software implementation risks will be introduced. The validity of GPT produced cryptographic function re-writes is highly testable for correctness and implementation flaws such as being incorrect for some proportion of inputs or memory leaks. A cryptographic algorithm serves as a good reference benchmark for the capabilities of generative machine learning to correctly re-write computer code because they need to implemented exactly correctly or they do not work – and in particular the code is therefore highly testable for correctness. Although there exist many instances of implementations of cryptographic algorithms in the codebases that very likely were used to train many of the GPT models in existence today, they are not as common (and in as many different varieties) as other foundational functions in computing (e.g., especially for teaching programming, such as sorting algorithms), and moreover their secure and correct implementation is extremely consequential.
@@ 2.1 GPT Model Implementation
The localGPT inference calls are performed on four Nvidia A100 GPU’s [9361255] with 82 Gigabytes of memory, with CUDA Version 12.4, and the GPT models are all obtained from the huggingface GPT repository [wolf2020huggingfaces]. The GPT models were trained and run using the Python 3 library PyTorch [Paszke_PyTorch_An_Imperative_2019].
@@ 2.2 GPT Output Parsing and Code Testing
    Count of how many of the function variants that were correct for compiler settings and had a Levenshtein character distance of 0 to the original source code (meaning, that the source code is strictly identical to this function variant), after repeated whitespace was removed and all comments were removed.
@@
    This step aggregates the compiler setting hash clustering into a unified graph (e.g., network) that reveals even more underlying clustering of the source code variants. The graph is defined by each binary being represented by a node, and edges are formed between nodes if the SHA-256 checksum of the two binaries is equal. Next, we check within all of these existing clusters if there exist any with hashes that are equal to the implementations with text distance of 0 to the original source code. This cluster is removed, and not included in the returned counts. In practice, we found there was always exactly 1 such cluster, and it was the largest cluster. The final metric that is reported is an aggregated number of how many disconnected components of this meta-graph exist, which corresponds to showing how many actually unique variants of this C code function were generated by the GPT models. Here we also report how many source code versions exist in each of the clusters of the meta-grouping. Note that function variants within each cluster may share identical syntax to each other – this clustering is specifically intended to delineate unique algorithmic invariant implementations of the original code. The integer counts of clustered group sizes are given in an unsorted sequence, separated by dashes.
@@
2.2.1 Correct Function Rewrite Composability
@@
Function Rewrites 30,322 36,343 40,422 30,379
Metric 1 10,680 7,349 18,517 10,416
@@
Metric $1\square$ 134 55 257 104
Metric 12* 48-23-17-18-17-18-54-22-17-17-21-17-22 116-53-38-39-38-39-118-55-56-56-59-56-61 121-76-73-74-71-74-149-74-69-69-70-69-70 67-29-18-24-20-24-90-19-17-17-19-17-19
@@ 3 Results
The aggregated test statistics defined in Section 2.2 are all reported, for each component function of the reference implementation, in Table 1. These metrics include a wide range of different inference temperature settings, the $1010$ different prompts, and are also aggregated from all $33$ GPT models. The most consequential metrics are the total number of attempted function re-writes, which is 137,466137466137,466137 , 466 in total, but only 46,9624696246,96246 , 962 of those could be compiled with all compiler settings. Fundamentally, this means that more than half of the GPT outputs could not be parsed as valid C code (using the markdown-style prompts as described in Sections 2.1 and 2.2). Note that the total number of function re-writes entry would be the exact same count (and exactly equal to the total number of GPT queries made) if not for the last parsing step (step number 7 in Section 2.2), where we occasionally get extra function re-writes from ancillary text produced by the GPT models.
@@
Across all source code variants and compiler settings, the binary execution fatal errors that were encountered were $93699369$ SIGSEGV errors, $562562$ SIGABRT errors, $99$ SIGILL errors, $33$ SIGFPE errors, and $1717$ SIGBUS errors. Across all source code variants and compiler settings $943943$ of the compiled binaries reached an apparent infinite loop state, as determined by a timeout check of 10 seconds.
The different hash string output lengths, as determined by Metric 40, across all of the function re-writes and compiled binaries are; 40,0,160,157286804001601572868040,0,160,1572868040 , 0 , 160 , 15728680. The vast majority of the generated hashes have length of $4040$, but one function re-write produces hashes (for some compiler optimization settings) of length $0$ (meaning an empty string), one case with a hash of length $160160$, and a different function re-write produces hashes of length $1572868015728680$ (for some compiler optimization settings), a majority of these characters are zeroes. Note that the testing code (Code Listing 12) is set up to read a specified number of indices from the array in which the hash data is computed.
@@
Figure 2 contain renderings of several of the connected components of the graphs produced by the correct SHA-1 re-write clustering, which are summarized by Metrics 12 and 13 in Table 1. Connections (e.g., paths and connected regions) of these graphs indicate that for some compiler optimization setting, an identical hash was produced for another piece of code (potentially the same code) with another compiler optimization setting. These connected graph components thus show function re-writes that are actually performing the same computation once compiled, even though the exact syntax of the code may be different. These graphs are all undirected. These graphs were generated using the Python 3 libraries Networkx [hagberg2008exploring] and Matplotlib [thomas_a_caswell_2021_5194481, Hunter:2007], and drawn using the kamada-kawai layout algorithm.
@@ 3.2 Example Function Variants and Compiled Binary Outputs
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ 4 Discussion and Conclusion
Fundamentally, using GPT models to produce source code, or as shown in this study to produce source code re-writes, is a cybersecurity risk for the integrity and stability of software development. The suite of tests that have been applied to these SHA-1 function re-writes has shown that GPT re-written functions can contain critical software flaws, some of which could be hard to detect without proper software validation and testing. GPT models have been shown to be incredibly effective at producing these SHA-1 re-writes at scale, but this capability of GPT models should serve as a research tool for testing interesting and useful versions of software, not as a solution for writing correct software. The tendency for GPT produced source code to contain bugs has been observed in several previous studies [sandoval2023lost, pearce2021asleep, spiess2024calibration, bhatt2023purple], meaning the findings of this study are consistent with the existing literature. The remarkable finding here is that these faulty code implementations were generated as a byproduct of GPT models with prompts requesting accurate code function re-writes, as opposed to other studies where the aim was to induce the generation of faulty code from GPT models [wu2023deceptprompt]. However, the other very notable finding of this study is the dramatic variety of code re-writes that were produced. While many of these re-writes were incorrect or contained serious software flaws, this suggests that GPT models offer a unique capability of producing large amounts of highly variable source code - which could be used for studying properties of computer code (namely, security properties), or fuzzing computer code.
@@
The prompting of GPT models unintentionally producing hash function implementations that fit the basic properties of hash functions, but are not correct to the actual algorithm (in this case SHA-1), is also very notable (Code Listings 7, 27, 28, 29, 30, 31, 32, 33 are explicit examples of these instances, but there are other instances of this occurring in other function re-write examples namely where the output is compiler optimization unstable). Fundamentally, being able to easily produce hash function implementations using generative transformer models that are valid hash functions (but not necessarily secure cryptographic hash functions), will make malware analysis more challenging. Not only does this change the source code, and therefore signatures of the compiled binaries, but it also makes automated cryptographic algorithm detection in binaries [7958617, 8960303, 6461007, 8392623, 10.1145/2382196.2382217, cryptoeprint:2022/1699, 5403858] significantly more challenging since these rely on finding known cryptographic constants used in standard cryptographic algorithms. Future work should analyze how secure these incorrect (e.g., non-SHA-1) hash function implementations are, for example using cryptographic fuzzing tools [10.1145/3628160, jin2022differential, ammann2024dy].
@@ Appendix C Checked Markdown Computer Code Language Identifiers
c, c++, Python, ruby, sql, java, go, css, perl, hpp, rust, php, md, markdown, ts, lua, bash, scss, csharp, kotlin, xml, matlab, vbnet, sh, yaml, vscode, arduino, objc, json, js, html, asp, console.log, text, txt, typescript, Makefile, asm, haskell, cpp, log, swift, #!lua, #!c, #!/bin/bash, #!/bin/sh, #!/usr/bin/perl, #! c++, #! sh
Note that, in many cases the markdown language identifier was not correctly identifying the language that was generated. Nevertheless, we do attempt to remove these language identifiers if they exist in order to increase the parse-ability of the GPT output.

### v2

@@ Abstract
Generative pre-trained transformers (GPT’s) are a type of large language machine learning model that are unusually adept at producing novel, and coherent, natural language. Notably, these technologies have also been extended to computer programming languages with great success. However, GPT model outputs in general are stochastic and not always correct. For programming languages, the exact specification of the computer code, syntactically and algorithmically, is strictly required in order to ensure the security of computing systems and applications. Therefore, using GPT models to generate computer code poses an important security risk – while at the same time allowing for potential innovation in how computer code is generated. In this study the ability of GPT models to generate novel and correct versions, and notably very insecure versions, of implementations of the cryptographic hash function SHA-1 is examined. The GPT models Llama-2-70b-chat-hf, Mistral-7B-Instruct-v0.1, and zephyr-7b-alpha are used. The GPT models are prompted to re-write each function using a modified version of the localGPT framework and langchain to provide word embedding context of the full source code and header files to the model, resulting in over 150,000150000150,000150 , 000 function re-write GPT output text blocks (that are potentially correct source code), approximately 50,0005000050,00050 , 000 of which were able to be parsed as C code and subsequently compiled. The generated code is analyzed for being compilable, correctness of the algorithm, memory leaks, compiler optimization stability, and character distance to the reference implementation. Remarkably, several generated function variants have a high implementation security risk of being correct for some test vectors, but incorrect for other test vectors. Additionally, many function implementations were not correct to the reference algorithm of SHA-1, but produced hashes that have some of the basic characteristics of hash functions. Many of the function re-writes contained serious flaws such as memory leaks, integer overflows, out of bounds accesses, use of uninitialised values, and compiler optimization instability. Compiler optimization settings and SHA-256 hash checksums of the compiled binaries are used to cluster implementations that are equivalent but may not have identical syntax - using this clustering over 100,000100000100,000100 , 000 distinct, novel, and correct versions of the SHA-1 codebase were generated where each component C function of the reference implementation is different from the original code.
@@ 1 Introduction
In this study we explore whether current GPT models can be used to generate correct algorithmic invariant implementations of a cryptographic hash function in C code. Specifically, we examine the task of rewriting an implementation of the cryptographic hash function known as SHA-1 [SHA1_NIST_specifications]. An interesting byproduct of code-rewriting with the GPT models is that a large number of the implementations are wrong in a variety of surprising ways, including containing software risks. These include compiler optimization instability (meaning that the output changes based on what compiler optimization settings were used), memory leaks, integer overflows, out of bounds writes, and implementations that are correct for some test vectors but not correct for other test vectors. Implementation risks in cryptographic algorithms is a critically important type of bug that exists in cryptography library implementations [cryptoeprint:2023/331, 9426079, 6494288, 8405614]. This study serves to caution that arbitrarily using GPT models for creating, or rewriting, source code can introduce serious flaws. GPT models, with their current capabilities, work well as research tools to study interesting source code variants, but using them for practical code generation poses a software security risk.
Importantly, SHA-1 [SHA1_NIST_specifications] is considered broken since at least one method of generating a SHA-1 hash collision has been demonstrated [stevens2017first], and is generally considered not secure for validating the integrity of data based on a variety of attacks [wang2005finding, biham2005collisions, karpman2015practical, stevens2013new, 255248]. However, SHA-1 is still widely deployed in information technology systems, and moreover its implementation is relatively short compared to other cryptographic algorithms, making it a good candidate for a proof of concept study on the capabilities of using GPT models for rewriting a cryptographic algorithm implementation.
@@
Automated code re-writing, in particular of cryptographic algorithm implementations, is an excellent test case for evaluating GPT models capabilities involving computer code because the implementation must be exactly correct or software implementation risks will be introduced. The validity of GPT produced cryptographic function re-writes is highly testable for correctness and implementation flaws such as being incorrect for some proportion of inputs or memory leaks. A cryptographic algorithm serves as a good reference benchmark for the capabilities of generative machine learning to correctly rewrite computer code because they need to implemented exactly correctly or they do not work – and in particular the code is therefore highly testable for correctness. Although there exist many instances of implementations of cryptographic algorithms in the codebases that very likely were used to train many of the GPT models in existence today, they are not as common (and in as many different varieties) as other foundational functions in computing (e.g., especially for teaching programming, such as sorting algorithms), and moreover their secure and correct implementation is extremely consequential.
@@ 2.1 GPT Model Implementation
The localGPT inference calls are performed on a small local cluster on four Nvidia A100 GPU’s [9361255] with 82 Gigabytes of memory, with CUDA Version 12.4, and the GPT models are all obtained from the huggingface GPT repository [wolf2020huggingfaces]. The GPT models were trained and run using the Python 3 library PyTorch [Paszke_PyTorch_An_Imperative_2019].
@@ 2.2 GPT Output Parsing and Code Testing
    Count of how many of the function variants that were correct for all compiler settings and had a Levenshtein character distance of 0 to the original source code (meaning, that the source code is strictly identical to this function variant), after repeated whitespace was removed and all comments were removed.
@@
    This step aggregates the compiler setting hash clustering into a unified graph (e.g., network) that reveals even more underlying clustering of the source code variants. The graph is defined by each binary being represented by a node, and edges are formed between nodes if the SHA-256 checksum of the two binaries is equal. Next, we check within all of these existing clusters if there exist any with hashes that are equal to the implementations with text distance of 0 to the original source code. This cluster is removed, and not included in the returned counts. In practice, we found there was always exactly 1 such cluster, and it was the largest cluster. The final metric that is reported is an aggregated number of how many disconnected components of the meta-graph exist (this combines binaries that were compiled with the same source code, in addition to the SHA-256 checksum formed graph), which corresponds to showing how many actually unique variants of this C code function were generated by the GPT models. Here we also report how many source code versions exist in each of the clusters of the meta-grouping. Note that function variants within each cluster may share identical syntax to each other – this clustering is specifically intended to delineate unique algorithmic invariant implementations of the original code. The integer counts of clustered group sizes are given in an unsorted sequence, separated by dashes.
@@
2.2.1 Composability of Correct Function Rewrites
@@
Function Rewrites 34,299 40,716 45,269 34,149
Metric 1 12,431 8,729 21,893 12,109
@@
Metric $1\square$ 133 75 398 111
Metric 12* 48-23-17-18-17-18-54-22-17-17-21-17-22 116-52-37-38-37-38-117-54-55-55-58-55-60 119-75-72-73-70-73-147-73-68-68-69-68-69 68-29-18-25-20-25-91-19-17-17-19-17-19
@@ 3 Results
The aggregated test statistics defined in Section 2.2 are all reported, for each component function of the reference implementation, in Table 1. These metrics include a wide range of different inference temperature settings, the $1010$ different prompts, and are also aggregated from all $33$ GPT models. The most consequential metrics are the total number of attempted function re-writes, which is 154,433154433154,433154 , 433 in total, but only 55,1625516255,16255 , 162 of those could be compiled with all compiler settings. Importantly, this means that more than half of the GPT outputs could not be parsed as valid C code (using the markdown-style prompts as described in Sections 2.1 and 2.2). Note that the total number of function re-writes entry would be the exact same count (and exactly equal to the total number of GPT queries made) if not for the last parsing step (step number 7 in Section 2.2), where we occasionally get extra function re-writes from ancillary text produced by the GPT models.
@@
Across all source code variants and compiler settings, the binary execution fatal errors that were encountered were $1068510685$ SIGSEGV errors, $559559$ SIGABRT errors, $99$ SIGILL errors, $33$ SIGFPE errors, and $1313$ SIGBUS errors. Across all source code variants and compiler settings $12631263$ of the compiled binaries reached an apparent infinite loop state, as determined by a timeout check of 10 seconds.
The different hash string output lengths, as determined by Metric 40, across all of the function re-writes and compiled binaries are; 40,0,160,2228264400160222826440,0,160,222826440 , 0 , 160 , 2228264. The vast majority of the generated hashes have length of $4040$, but one function re-write produces hashes (for some compiler optimization settings) of length $0$ (meaning an empty string), one function re-write produces a hash of length $160160$, and a different function re-write produces hashes of length $22282642228264$ (for some compiler optimization settings), a majority of these characters are zeroes. Note that the testing code (Code Listing 12) is set up to read a specified number of indices from the array in which the hash data is computed.
@@
Figure 5 contain renderings of several of the connected components of the graphs produced by the correct SHA-1 re-write clustering, which are summarized by Metrics 12 and 13 in Table 1. Connections (e.g., paths and connected regions) of these graphs indicate that for some compiler optimization setting, an identical hash was produced for another piece of code (potentially the same code) with another compiler optimization setting. These connected graph components thus show function re-writes that are actually performing the same computation once compiled, even though the exact syntax of the code may be different. These graphs are all undirected. These graphs were generated using the Python 3 libraries Networkx [hagberg2008exploring] and Matplotlib [thomas_a_caswell_2021_5194481, Hunter:2007], and drawn using the kamada-kawai layout algorithm. Note that, counter-intuitively, more SHA-1 rewrites can actually cause more connections to be revelealed by the compilers, and thus fewer unique function clusters be found by Metric 13.
Figure 1: Correct code rewrite performance metrics as a function of inference temperature.
Figure 2: Correct code rewrite performance metrics as a function of the $1010$ different prompts.
Figure 3: Correct code rewrite performance metrics as a function of the three GPT models.
So as to compare the various inference temperatures, GPT models used, and prompts, we use three high level metrics (that do not depend on the clustering of the various versions produced by different compiler settings); Metric 1 (was able to be compiled for all compiler settings), Metric 3 (compiled binary was output-verified for all compiler settings), Metric 7 (the compiled binary was output-verified for all compiler settings, and had source code distance greater than 0 compared to the original) as outlined in Section 2.2. For all three of these summary metrics, the larger they are the better the model and parameter choice is performing. Metric 7 is the most fundamental and important measure of code quality, and also being measurably different from the original source code. Figure 1 shows these three metrics as a function of inference temperature (across all $1010$ prompts and $33$ GPT models). Figure 2 shows these three metrics as a function of the $1010$ different prompts. Figure 3 shows these three metrics as a function of the $33$ different GPT models.
There are several clear findings from these aggregate performance metrics. Figure 1 shows that the proportion of compilable functions increases consistently as the temperature gets lower. However, Metric 7 peaks at a temperature of $0.40.4$ to $0.20.2$ - this shows that some stochasticity in the token choice is required for the generated code to be different from the original code, but too high stochasticity makes the code quality decrease as well. Figure 2 shows that the best performing prompt to maximize Metric 7 was Prompt number 7, and the worst prompt to maximize Metric 7 was Prompt number 5. Finally, Figure 3 shows that Llama-2-70b-chat-hf performed the best.
@@ 4 Discussion and Conclusion
Fundamentally, using GPT models to produce source code, or as shown in this study to produce source code re-writes, is a cybersecurity risk for the integrity and stability of software development. The suite of tests that have been applied to these SHA-1 function re-writes has shown that GPT re-written functions can contain critical software flaws, some of which could be hard to detect without proper software validation and testing. GPT models have been shown to be incredibly effective at producing these SHA-1 re-writes at scale, but this capability of GPT models should serve as a research tool for testing interesting and useful versions of software, not as a solution for writing correct software. The tendency for GPT produced source code to contain bugs has been observed in several previous studies [sandoval2023lost, pearce2021asleep, spiess2024calibration, bhatt2023purple], meaning the findings of this study are consistent with the existing literature. The remarkable finding here is that these faulty code implementations were generated as a byproduct of GPT models with prompts requesting accurate code function re-writes, as opposed to other studies where the aim was to induce the generation of faulty code from GPT models [wu2023deceptprompt]. However, the other very notable finding of this study is the dramatic variety of code re-writes that were produced. While many of these re-writes were incorrect or contained serious software flaws, this suggests that GPT models offer a unique capability of producing large amounts of highly variable source code - which could be used for studying properties of computer code (namely, security properties), or fuzzing computer code. The code re-writing procedure described in this study should also be applied to implementations of other cryptographic algorithms to determine how well the GPT models perform at rewriting potentially even more complex code than this SHA-1 implementation.
@@
The prompting of GPT models unintentionally producing hash function implementations that fit the basic properties of hash functions, but are not correct to the actual algorithm (in this case SHA-1), is also very notable (Code Listings 7, 27, 28, 29, 30, 31, 32, 33 are explicit examples of these instances, but there are other instances of this occurring in other function re-write examples namely where the output is compiler optimization unstable). Fundamentally, being able to easily produce hash function implementations using generative transformer models that are valid hash functions (but not necessarily secure cryptographic hash functions), will make malware analysis more challenging. Not only does this change the source code, and therefore signatures of the compiled binaries, but it also makes automated cryptographic algorithm detection in binaries [7958617, 8960303, 6461007, 8392623, 10.1145/2382196.2382217, cryptoeprint:2022/1699, 5403858] significantly more challenging since these rely on finding known cryptographic constants used in standard cryptographic algorithms. Future work should analyze how secure these incorrect (e.g., non-SHA-1) hash function implementations are, for example using cryptographic fuzzing tools [10.1145/3628160, jin2022differential, ammann2024dy]. More detailed analysis of the GPT rewritten code, specifically the compiled binaries, can be performed in future research - namely things such as memory usage, compute time efficiency, and detecting other potential security flaws such as side channel attack susceptibility.
@@ Appendix C Checked Markdown Computer Code Language Identifiers
List of markdown and miscellaneous code identifiers that are checked for after the triple back-tick output. This list is delineated by commas.
c, c++, python, ruby, sql, java, go, css, perl, hpp, rust, php, md, markdown, ts, lua, bash, scss, csharp, kotlin, xml, matlab, vbnet, sh, yaml, vscode, arduino, objc, json, js, html, asp, console.log, text, txt, typescript, makefile, asm, haskell, cpp, log, swift, #!lua, #!c, #!/bin/bash, #!/bin/sh, #!/usr/bin/perl, #! c++, #! sh, javascript, json..., assembly, #!, awk, [c], temp, triple backquotes, uint32_t, [triple backquotes], uint8_t, document, backquote, instructions, bitlen, uint, word, byte, size_t, const, fsharp, csh, typename, std::size_t, state[], xxd, "hello, world!", data[], data[ ], csv, shellscript, erb, plaintext, ocl, shell
Note that, in nearly all of these cases the markdown language identifier was not correctly identifying the language that was generated, and actually include unrelated utility Linux commands or general programming keywords. Nevertheless, we do attempt to remove these language identifiers if they exist in order to increase the parse-ability of the GPT output. Each of these markdown identifiers is checked against the first line of strings in the parsed output, with all capitalization having been removed. Additionally, cases such as miscellaneous spaces before the markdown identifiers are checked for, as well as extra ancillary backticks; if any of these cases are found, the ancillary is removed.